### v1

# Imaging supermoiré relaxation and conductive domain walls in helical trilayer graphene

Abstract
In twisted van der Waals materials, local atomic relaxation can significantly alter the underlying electronic structure and properties. Characterizing the lattice reconstruction and the impact of strain is essential to better understand and harness the resulting emergent electronic states. Here, we use a scanning single-electron transistor to image spatial modulations in the electronic structure of helical trilayer graphene, demonstrating relaxation into a superstructure of large domains with uniform moiré periodicity. We further show that the supermoiré domain size is enhanced by strain and can even be altered in subsequent measurements of the same device, while nevertheless maintaining the same local electronic properties within each domain. Finally, we observe higher conductance at the boundaries between domains, consistent with the prediction that they host counter-propagating topological edge modes. Our work confirms that lattice relaxation can produce moiré-periodic order in twisted multilayers, demonstrates strain-engineering as a viable path for designing topological networks at the supermoiré scale, and paves the way to direct imaging of correlation-driven topological phases and boundary modes.
I Introduction
The electronic properties of quantum materials depend sensitively on their lattice structure. Compared to conventional solids, the weak interlayer bonding in van der Waals systems provides additional freedom for atomic-scale relaxation driven by the competing energetics of different stacking configurations and elastic strain [1, 2, 3, 4]. As a result, local twists and strains can develop, enlarging areas with preferred stacking while shrinking disfavored arrangements [5, 6, 7, 8, 9, 10, 11]. This has profound effects on both band structure and the resulting electronic states. For example, different interlayer shifts produce distinct electronic structure [12, 13, 10, 14, 15], and atomic displacements can even alter band topology [16]. Lattice relaxation can also generate new functionality, ranging from alternating ferroelectric domains driven by local symmetry breaking [17, 18, 19, 20, 21] to solitons which host topologically protected boundary modes [22, 23, 24] and can form networks of one-dimensional edge states [25, 5, 6, 7, 26, 27, 8, 9]. Beyond fundamental interest, designing systems where individual edge modes can be controlled may enable applications in topology-based nanoelectronics [22, 23, 25, 28].
Helical trilayer graphene (HTG), in which three graphene layers are consecutively twisted by equal angles , provides an especially rich venue to explore the relationship between lattice relaxation and electronic structure [29, 30, 31, 32, 33, 34, 4, 35, 36, 37, 38, 39]. Nominally, HTG is quasi-periodic, as the moiré pattern formed by the first and second layers is incommensurate with that formed by the second and third layers. However, at small twist angles, HTG is predicted to relax into large triangular domains of moiré-periodic order [32, 33, 34, 4, 35]. This yields two key length scales: a moiré wavelength , which is uniform within the individual domains; and a larger supermoiré wavelength of order , which characterizes the size of the domains. Here is the lattice constant of graphene. HTG has two inequivalent types of domains (denoted as and in Fig. 1a) with different atomic stacking configurations. The domains locally break symmetry, but map onto each other under a transformation (Fig. ED1). Both domains locally host low energy bands that are separated from more dispersive bands by a large band gap.
Crucially, the Dirac cones of the three graphene layers, together with the superlattice gap induced by locally broken symmetry, endows the low-energy bands of each valley with non-zero Chern numbers. Adjacent domains have opposite Chern number for a given valley such that HTG realizes a network of topological boundary modes when the domains are tuned to incompressible fillings [32, 33]. Furthermore, near a magic angle , the topological bands become flat, making HTG a promising platform to investigate interaction-driven quantum anomalous Hall states [32, 33, 34, 38, 39]. Indeed, a recent transport study of magic-angle HTG observed robust, but non-quantized anomalous Hall signals at both integer and fractional filling factors [36]. It remains unknown whether the absence of quantization is related to the presence of domain walls. The above theoretical expectations and experimental signatures strongly motivate direct characterization of the lattice relaxation, its effect on electronic structure, and the role of domain walls, none of which have been experimentally addressed to date.
Here, we conduct single-electron transistor (SET) microscopy of HTG, which reveals periodic modulations of the local electronic structure at the supermoiré scale. Our measurements indicate that HTG relaxes into large domains of moiré-periodic order that locally break symmetry. We demonstrate decoupling between moiré and supermoiré length scales, with strain capable of enlarging and reshaping domains without changing the local electronic character within them. We also observe increased conductivity along the network of domain walls when the domains are gapped, consistent with the prediction of helical edge modes. Our work clarifies the interplay between lattice reconstruction and strain, provides key insight for interpreting transport measurements of mesoscopic devices, and establishes HTG as a promising platform for tunable topological networks.
II Electronic structure and lattice relaxation in HTG
A schematic of the scanning SET measurement setup is illustrated in Fig. 1a (see Methods), and an optical micrograph of the HTG device is shown in Fig. 1b. We plot the inverse electronic compressibility d/d as a function of carrier density at a representative location within the sample in Fig. 1c (red curve). A pair of prominent incompressible peaks occur near cm-2, which we attribute to the superlattice gaps at moiré filling factors . These gaps separate the flat bands near the charge neutrality point from the remote dispersive bands (Fig. 1d). From the carrier densities of the superlattice peaks, we infer a local twist angle (Methods), and as we discuss below, the observation of only one pair of superlattice peaks indicates a single uniform moiré periodicity, despite the presence of two interlayer twist angles. The change in chemical potential across the flat bands is extremely small, meV (Fig. 1c, blue curve). The above characteristics match well to theoretical calculations of the HTG band structure [32, 33, 34, 35, 38].
We also observe a sharp incompressible peak at charge neutrality accompanied by pronounced electron-hole asymmetry. This is at odds with prior theoretical models, which predict gapless Dirac nodes at the charge neutrality point and nearly electron-hole symmetric bands [32, 33]. The discrepancy can be reconciled by including momentum-dependent tunneling terms in the band structure calculation, which open a small band gap and qualitatively reproduce the asymmetric behavior (Supplementary Sec. 1), underscoring their importance for accurate modeling of the electronic structure. We also comment that while is smaller than that in magic-angle twisted bilayer [40, 41] and trilayer graphene [13, 42], we do not observe correlation-driven gaps at intermediate filling factors. This suggests that the underlying single-particle bandwidth is a better proxy for strong interaction effects than , which can be renormalized (broadened) by interactions [43, 44].
A spatial line cut along the black trajectory in Fig. 1b provides the first indication of lattice relaxation in HTG. The electronic character is uniform (Fig. 1e), and the superlattice peaks disperse only weakly, indicative of small local twist angle variations (Fig. 1f). Importantly, we observe only a single pair of superlattice peaks throughout the spatial line cut. The same observation holds for the entire area of the device (Fig. ED2). This is notable because HTG has two interlayer angles, moiré systems are notoriously plagued by uncontrolled twist angle disorder [45, 46], and spatial variations in each interlayer angle need not be correlated. One would then generically expect two distinct interlayer twist angles in a given location and thus two pairs of superlattice peaks, as observed in other twisted trilayer graphene systems [47, 48]. However, based on the width in carrier density of the superlattice peaks, the interlayer angles would have to match to within throughout the device to produce a single pair of peaks, which we view as extremely unlikely. We also rule out that a large angle mismatch leads to a second superlattice peak outside the accessible range of carrier densities (Supplementary Sec. 2). Instead, we interpret the single pair of peaks as evidence for local relaxation that produces moiré-periodic order within the domains. Theory predicts that this is favored even when starting from two slightly mismatched twist angles, in which case the moiré periodicity is determined by an average of the interlayer angles [32, 33, 34, 4].
III Imaging supermoiré domains
To further demonstrate the effects of lattice relaxation in HTG, we measure the spatial dependence of d/d in the vicnity of the superlattice peak within a 10 area (white box in Fig. 1b). The local twist angle (Fig. 2a) is remarkably homogeneous. In contrast, the peak value of d/d at each location (Fig. 2b) exhibits approximately periodic spatial modulations that are not correlated with features in the twist angle map. The modulations occur on a length scale of a few hundreds of nanometers, significantly exceeding that of the moiré wavelength nm. We therefore ascribe them to variations of the electronic bands in HTG at the supermoiré scale. Similar modulations with the same periodicity are also observed for all other measured incompressible states, both at zero magnetic field (Fig. ED3) and in a perpendicular magnetic field (Fig. ED4).
The locations where d/d at has local minima (dark spots, Fig. 2b) form a triangular lattice. Surrounding these sites, the local maxima form a honeycomb structure (green, Fig. 2b), with narrow lines of lower d/d separating adjacent maxima. This broadly agrees with the theoretical prediction that HTG relaxes into triangular and domains, separated by domain walls where the local stacking order switches (Fig. 1a, right) [32, 33, 34, 4, 35]. In particular, the triangular domains are predicted to be fully gapped at , while their boundaries host gapless modes (see further discussion below). As a result, d/d is maximized at the domain centers, while the gapless modes lead to enhanced compressibility at the domain walls, reaching a local minimum in d/d where they meet at AAA stacking sites (Supplementary Sec. 1). We therefore associate the local minima in Fig. 2b with AAA stacking sites, and the local maxima with the centers of the and domains.
Building on this qualitative understanding, we next assess the length scale of the supermoiré modulation more quantitatively. The experimentally observed distance between adjacent AAA sites is of the same order of magnitude as the supermoiré wavelength nm expected from the local twist angle of the device. However, variations in AAA site separation produce a range of supermoiré domain areas, defined as the triangular areas enclosed by adjacent AAA sites (Fig. 2c). We compare the experimentally observed areas to the theoretically predicted value assuming equal interlayer angles and zero strain. Their ratio is encoded in the color of each triangle in Fig. 2c, which reveals substantial enhancement in the supermoiré domain area relative to the theoretical prediction over a large subset of the device. We emphasize that this cannot be explained by twist angle mismatch because monotonically decreases with increasing (Fig. 2g, Supplementary Sec. 2).
We instead attribute the enhanced supermoiré areas to strain, which has been proposed to modulate domain size in HTG [32] and is ubiquitous in moiré systems [7, 8, 9, 10, 46, 49]. For simplicity, we first discuss isotropic biaxial heterostrain applied to the middle graphene layer, where denotes stretching. The supermoiré wavelength is enhanced within a narrow range centered around a divergence at (Fig. 2h). Strain therefore provides a natural explanation for the increased domain sizes. We note that other strain configurations can also enlarge domains, and uniaxial strain is required to account for the anisotropy of the observed supermoiré domains. We provide a more comprehensive discussion of other strain configurations in Supplementary Sec. 3.
Remarkably, we discover a dramatic reshaping of the supermoiré domains in subsequent measurements of the same device after thermal cycling and other changes (see Supplementary Sec. 4 for details). A second round of measurements in the area indicated by the red box in Fig. 1b (largely overlapping with the bottom part of Fig. 2a-b) is shown in Fig. 2d-e. While the local twist angle is nearly identical, the supermoiré domain sizes substantially increase, up to 4 times the theoretically predicted area (Fig. 2f, i). Moreover, the supermoiré domains are more isotropic relative to those observed in the first round of measurements.
The areal enhancement of the domains, as well as the change in shape, can be attributed to redistribution of strain between measurements. In the new configuration, larger domains are favored to minimize regions associated with higher stacking energy, i.e. domain walls and AAA sites [1, 3, 4, 32], with a relatively smaller cost of increased elastic strain energy. We emphasize that this reshaping of supermoiré domains occurs with almost no change in the local twist angle (i.e. moiré periodicity) or the band structure within the domains, as discussed in more detail at the end of the manuscript. This demonstrates a separation of length scales, providing new opportunities for strain engineering of supermoiré-scale networks without perturbing local physics.
IV Enhanced domain wall conductance
We next address the electronic character of the domain walls in HTG, utilizing simultaneous, but independent measurements on a.c. and d.c. timescales with the SET (Methods). High-resolution spatial maps of the step in chemical potential at measured with each modality (Fig. 3a-b) show qualitatively similar supermoiré modulations to those in Fig. 2b: is maximized at the domain centers and minimized at the AAA sites. Quantitatively however, is larger than , and it exhibits higher contrast. While both measurements should produce identical results when the sample conducts well, the a.c. modality can exhibit spurious enhancement if sample resistance becomes large enough to make the time constant comparable to the inverse a.c. measurement frequency (Supplementary Sec. 5). Therefore, by studying local variations in the degree of a.c. enhancement, we can qualitatively characterize the relative local conductivity of HTG.
The degree of a.c. enhancement, defined in Fig. 3c, exhibits systematic spatial dependence across the supermoiré domain structure (Fig. 3d). It is largest in the center of the and domains, is reduced at the domain walls, and is smallest at AAA sites. This indicates that the boundaries between domains are more conductive than their interiors (Supplementary Sec. 5). Theoretically, counter-propagating topological edge states with opposite Chern number for the two valleys are predicted along domain walls at in HTG (Fig. 3e). These edge modes manifest as remote Dirac cones within the superlattice gaps of local band structure calculations at the domain wall and AAA site (Fig. 3f-g). While we only probe relative conductivity and thus do not directly address topology, our results are consistent with theoretical predictions.
To directly compare the data and theory, we calculate from the HTG band structure in domains, at domain walls, and at AAA sites, taking and accounting for the finite displacement field V/nm at (Supplementary Sec. 1). A spatial map of the theoretically calculated is shown in Fig. 3h. After accounting for the finite resolution of the SET, the simulated spatial dependence of produces a pattern that matches well with experiment (Fig. 3i, Supplementary Sec. 6).
V Magnetic field dependence
Finally, we discuss the behavior of HTG in a perpendicular magnetic field . A Landau fan of d/d as a function of and (Fig. 4a) reveals an intricate pattern of quantum Hall and Hofstadter states. Each can be classified by an integer Chern number and zero-field intercept . Within different regions of the Landau fan, we observe states that extrapolate back to every integer filling factor between , as shown in the Wannier diagram of Fig. 4b.
Qualitatively, the Landau fan is consistent with the low-energy band structure of HTG discussed above and can largely be captured by single-particle Hofstadter spectrum calculations (Fig. 4c, Supplementary Sec. 1). We observe an electron-hole asymmetry due to momentum-dependent tunneling, and the most prominent states emanating from the charge neutrality point with are consistent with the presence of two (weakly gapped) Dirac cones at the and points (Fig. 1d).
Multiple features in the Landau fan also indicate symmetry breaking, which leads to distinct Hofstadter spectra for the and valleys for . Specifically, the absence of large Hofstadter gaps at high fields near the charge neutrality point can be traced to crossing Hofstadter subbands that originate from different valleys (Fig. ED6, Supplementary Sec. 1). Similarly, the presence of Hofstadter states that extrapolate to intermediate integer are a consequence of the different Hofstadter spectra for each valley. We comment that these can arise without interaction effects. However, the weaker states with odd-integer Chern numbers or zero-field intercepts require spin splitting and likely reflect Hofstadter subband ferromagnetism [50, 51, 52].
In Fig. 4a, we observe neither evidence of moiré-of-moiré physics with states emerging from incommensurate filling factors [53], nor evidence for the presence of two distinct moiré periodicities [47, 48]. This further supports the conclusion that HTG relaxes into large moiré-periodic domains. Moreover, the Landau fan in Fig. 4a is generic across the entire sample (Fig. ED7) and the pattern of Hofstadter states was virtually identical in the second round of measurement (Fig. ED8), despite the significant rearrangement of the supermoiré domain shapes and sizes. These findings emphasize that although heterostrain can modulate the supermoiré pattern, the local electronic properties remain unchanged.
VI Conclusion
In conclusion, we directly imaged spatially modulated electronic states in HTG at supermoiré length scales, demonstrating a profound influence of lattice relaxation on electronic structure. The existence of moiré-periodic domains with conductive edge states at their boundaries has important implications for interpreting mesoscopic transport measurements. Our work strongly motivates future imaging and spectroscopic efforts to study the interactions between edge states and the degree to which they are protected from backscattering [6, 54]. The separation of the moiré and supermoiré length scales indicates that strain can be used to design topological networks, as well as obtain single-domain devices. Due to the ideal quantum geometry of HTG at the magic angle, this may enable the realization of strongly-correlated topological states with quantized response [32, 55, 56].